OpenStudy (mallorysipp234):

I have a test tomorrow and i was wondering if theres anyway possible that i can divide a polynomial into something like 2x-2 in the calculator itself WITHOUT doing the long division. I dont wanna fail cuz of ONE problem

Nnesha (nnesha):

are you familiar with synthetic division? that's like addition/subtraction

OpenStudy (mathmale):

Depends upon which calc you have. There was a TI-92 twenty years ago that could factor in some situations. Is factoring your main goal? If so, consider Nnesha's suggestion of using synthetic division to find the roots of given polynomials. From these roots it's a snap to write out the corresponding factors.

Nnesha (nnesha):

btw long division is really easy if you understand the steps

OpenStudy (mallorysipp234):

I am familiar with synthetic division and it makes a lot more sense to me than the regular "long division" process. I dont know how to do it with something like "2x-2" i CAN however do a synthetic with something like "x-4" its too much room for error and confusion for me.

OpenStudy (shadowlegendx):

Dude, I remember when I thought division with polynomials was hard, but then I learned it and aced that quiz :)

OpenStudy (mathmale):

Perhaps you could post a polynomial or two and see what we can do with synthetic div in finding roots and then factors. I dont know how to do it with something like "2x-2" .... In your shoes I'd factor out the 2, obtaining (x-2), and then use the latter as the divisor in long div, or use 2 as the divisor in synth div. Try to narrow down what you're asking here. How can we help you at this moment?

OpenStudy (shadowlegendx):

Long Division is super easy

OpenStudy (mathmale):

Personally I prefer synth. div.; it's sometimes just a matter of preference which one u use.

OpenStudy (mallorysipp234):

so per say, can i use a calculator to divide (8x^3+18x^2+18x+11) / (2x+2)?????????????

OpenStudy (mallorysipp234):

without the awful steps in a long division process. Im timed on my test and it would take me forever.

OpenStudy (mathmale):

Again, Mallory, whether or not your calculator would help depends on which one you have. I'm going to try synth. div. on your example, quickly, and see if that helps me to help u.

OpenStudy (mallorysipp234):

I have another more simple question absolutely separate from this after and ok thank u!

Nnesha (nnesha):

1) divide first terms \[\large \rm \frac{ first ~term~of~polynomail }{ first~term~of~divisor }\] 2: multiply(distribute) 3: subtract all terms 4: carry down 5: repeat  \[\rm \frac{ 8x^3}{ 2x} =2x^2\] and then distribute divisor by 2x^2